Nonconstant growth Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $0.75 coming 3 years from today. The dividend should grow rapidly - at a rate of 48% per year - during Years 4 and 5; but after Year 5, growth should be a constant 6% per year. If the required return on Computech is 13%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations. $
Answer #1

D3=0.75

D4=(0.75*1.48)=1.11

D5=(1.11*1.48)=1.6428

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(1.6428*1.06)/(0.13-0.06)

=24.8766857

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=0.75/1.13^3+1.11/1.13^4+1.6428/1.13^5+24.8766857/1.13^5

=$15.59(Approx).